# Attribute First, then Generate: Locally-attributable Grounded Text Generation

### v1

@@ Abstract
Recent efforts to address hallucinations in Large Language Models (LLMs) have focused on attributed text generation, which supplements generated texts with citations of supporting sources for post-generation fact-checking and corrections. Yet, these citations often point to entire documents or paragraphs, burdening users with extensive verification work. In this paper, we introduce a locally-attributable text generation approach, prioritizing concise attributions. Our method, named “Attribute First, then Generate”, breaks down the conventional end-to-end generation process into three intuitive steps: content selection, sentence planning, and sequential sentence generation. By initially identifying relevant source segments (“select first”) and then conditioning the generation process on them (“then generate”), we ensure these segments also act as the output’s fine-grained attributions (“select” becomes “attribute”). Tested on Multi-document Summarization and Long-form Question-answering, our method not only yields more concise citations than the baselines but also maintains—and in some cases enhances—both generation quality and attribution accuracy. Furthermore, it significantly reduces the time required for fact verification by human assessors.
@@
Figure 1: In our approach, demonstared here on LFQA, the input to the model assumes a question and a collection of documents, which could be retrieved or provided by the user. The generated text is grounded in these documents, accompanied with sentence-level citations to concise relevant spans from the supporting sources. Previous approaches (e.g. Gao et al., 2023a, b) cite the whole source (truncated here for brevity), making it harder for readers to reliably verify the generation.
@@ 3.2 In-context Learning
For each of the substeps delineated in the previous section, we propose a prompt-based in-context learning (ICL) strategy, where a model is guided with specific instructions and carefully selected few-shot examples. The content selection step ((a) in Fig. 2) is formulated as $\mathcal{D}\rightarrow h_{1},\ldots,h_{m}$, where $\mathcal{D}$ is the set of source documents and $h_{i}$ corresponds to a verbatim span from $\mathcal{D}$. Namely, a model is tasked with outputting relevant spans from $\mathcal{D}$, separated by a designated markup (see prompt in Fig. B.1). In MDS, these spans are aimed to capture salient information, while for LFQA, the model, presented with a query, is instructed to identify spans answering that query. These spans are then located within the source documents via string-matching, with any undetectable span being omitted (for more details, see Section C.1). Notably, in all subsequent steps, we “highlight” the identified spans within the document set $\mathcal{D}$ via designated markups before and after each span, akin to the methodology used in prior work (Gehrmann et al., 2018; Slobodkin et al., 2022, 2024).
For the sentence planning step ((b) in Fig. 2), the documents, now “highlighted”, are concatenated, truncating each at the final highlighted span to manage input length and focus the model’s attention. The model is prompted to organize $h_{1},\ldots,h_{m}$ into an ordered set of highlight clusters $C_{1},\ldots,C_{n}$, where $C_{i}=\{h_{k}\}_{k=1}^{|C_{i}|}$, to guide the subsequent generation phase (see prompt in Fig. B.2). For more details, see Section C.2.
Finally, in step (c), the model is guided to generate the output one sentence at a time, adhering to the structured plan. Specifically, we separate the generation process into a series of model calls, each focusing on a single cluster’s highlights, concatenated in the same manner as in step (b). The model is also presented with all previously generated sentences $s_{1:i}$. The task is then structured as maximizing $p(s_{i+1}|s_{1:i},C_{i+1})$, guiding the model to produce the subsequent sentence $s_{i+1}$, considering both the current cluster of highlights $C_{i+1}$ and the previously-generated sentences $s_{1:i}$. [^1] In Section 5.1 we show that conditioning on $s_{1:i}$ improves coherence. For more details, see Section C.3 as well as the full prompt structure in Fig. B.3.
@@
We also explore a variation that combines the sentence-planning and sentence-by-sentence generation stages ((b) and (c)), inspired by the Chain-of-Thought paradigm (CoT; Wei et al., 2022). Given the selected content from step (a), the model is tasked with generating sentence-level plans, deciding on highlights to include in each synthesized sentence, alongside producing these sentences (see prompt in Fig. B.4). The model is guided to produce the final output, after forming the plan, focusing on unifying and refining sentences for coherence. For sentence-level attribution, we link each sentence in the output with the highlights from its planning stage in the CoT process. [^2] We use spaCy sentencizer Honnibal and Montani (2017). For more details, refer to Section C.3.
@@ 3.3 Fine-tuning
We also explore fine-tuning models for the various steps of our framework, aligning the input-ouput formulation of each subtask to its ICL counterpart in Section 3.2. During content selection ((a) in Fig. 2), our model is taught to copy task-specific segments from $\mathcal{D}$ verbatim, and to separate them using a special markup (see Section C.1 and Fig. B.5). To further ensure strict lexical fidelity to the source text, a constrained decoding approach is adopted during inference (see Section C.6 for more details).
In the sentence planning step ((b) in Fig. 2), we use special tokens to ’highlight’ the previously selected spans, and teach our model to copy verbatim the “highlighted” spans in a clustered and ordered manner, using special tokens to separate between different clusters and different highlights within a cluster (see Section C.2 and Fig. B.6 for more details). This phase similarly incorporates a constrained decoding technique to maintain close lexical similarity to the highlighted content. Finally, for the sentence generation step ((c) in Fig. 2), we use special tokens to separate between the highlighted documents, the prefix, and for the LFQA setting the query (see Section C.3 and Fig. B.7).
@@
We also explore a joint strategy that combines content selection and plan generation. Here, the entire document set $\mathcal{D}$ serves as input, and the output is an ordered series of “highlight” clusters $C_{1},\ldots,C_{n}$. To ensure exact copying of source spans, we employ the aforementioned constrained decoding strategy. For more details, see Section C.2.
@@ 4.1 Datasets
For MDS, we use an alignment dataset derived from the DUC, TAC benchmarks [^3] https://duc.nist.gov/ and the Multi-News dataset (Fabbri et al., 2019), comprising of pairs of summaries, corresponding source articles and summary-source alignments. [^4] https://anonymous.4open.science/r/multi-news-dataset-5B4F and https://github.com/oriern/SuperPAL. For LFQA, we rely on human-annotated alignments of large language model responses with citations Liu et al. (2023). By filtering responses that were rated as verifiable, we ensure high-quality alignment for this setting. Further details and dataset processing information are provided in Appendix A.
@@ 4.2 Models and Baselines
Table 1: MDS results of both ICL and fine-tuned models. The left two columns are metrics that measure the quality of the output, ignoring attributions. The right columns report the attribution quality, including the length (number of tokens) of the cited text. No Att. shows the percent of generated sentences lacking any attribution.
@@
Table 2: LFQA results of both ICL and fine-tuned models. The left two columns are metrics that measure the quality of the output, ignoring attributions. The right columns report the attribution quality, including the length (number of tokens) of the cited text. No Att. shows the percent of generated sentences lacking any attribution.
@@ 4.3 Evaluation
The evaluation of generated attributed texts consists of two steps. First, annotators are asked to evaluate text coherence on a 5-point Likert scale, and in the LFQA setting also its helpfulness in addressing the query. Next, annotators proceed to evaluate the quality of citations. We follow the sentence-level attribution evaluation protocol from Gao et al. (2023a), which extends the Attributable to Identified Sources (AIS; Rashkin et al., 2022) human evaluation framework of attribution to the sentence-level. Annotators judge whether each output sentence is fully supported by its attribution, scoring 1 for full support and 0 otherwise. The final AIS score is the average of these scores across sentences. For more details, see Appendix F.
@@ 5 Results and Analyses
We first note that our approach, which extracts focused spans from the input documents, leads to significantly shorter attributions than the ALCE baseline, where citations coarsely point to entire documents. For example, the Attr. FirstCoT variant’s citations are on average 45 times shorter than ALCE’s. Importantly, we find that assessing our paradigm’s attribution takes annotators roughly half the time compared to those of ALCE, emphasizing the benefit of attribution localization in streamlining the manual fact-checking process.
In terms of attribution accuracy, our method surpasses the ALCE baseline in the LFQA domain (both on the automatic AutoAIS and the human AIS metrics), while in MDS, although ALCE shows higher accuracy, it leads to twice as many unattributed sentences. We also note that while in the MDS setting our method may decrease attribution accuracy compared to ALCE’s full-document citations, readers can always fall back to examining the full document if the provided highlights are insufficiently supporting. Indeed, analyses of the human annotations revealed that around 42% of the 109 sentences deemed unsupported were partially supported, with the remainder of the attributing details frequently found in immediate proximity to the highlighted sections (usually within the same or adjacent sentences). [^8] See partially-attributed examples in Appendix G. This observation indicates that while our sentence fusion step occasionally includes slightly more information than the selected highlights (later used as fine attributions), it mostly adds details from the surrounding context. [^9] This trend is also observed in our automatic evaluation, where extending sub-sentence attributing spans to their corresponding full sentences improves AutoAIS. See Appendix D. Hence, while suggesting room for improvement in strict adherence to highlights, this does not significantly hinder fact-checking effectiveness.
@@ 5.1 Do Models Benefit from Seeing the Prefix?
We also examine whether providing models with the context of previously generated sentences indeed aids the generation process, particularly in enhancing output cohesion as defined by Maimon and Tsarfaty (2023). Cohesion involves sentences being “referentially linked”, e.g., via co-referring mentions, or linked through discourse connectors. To analyze the impact of the prefix on cohesion, we modified the sentence generation prompt to exclude the prefix and conducted the sentence fusion step (step (c) in Fig. 2) using the ICL variant for both MDS and LFQA tasks, maintaining identical sentence-level plans. Starting from the second sentence to isolate the impact of the prefix, we examined 20 instances per task. Our analysis of 113 sentences reveals that approximately 30% were more cohesive when including the prefix. Specifically, 23% showed improved referential linkage, mostly attributed to better use of co-reference, and 7% demonstrated better discourse connector usage. [^10] See Appendix H for examples. Additionally, our findings suggest that prefixes play an increasing role in cohesion as the output progresses (see Fig. 3), implying that cohesion gains importance as the text develops. [^11] The observed plateau in cohesion levels in LFQA is attributed to its outputs generally not exceeding three sentences.
@@ 7 Conclusion
Our Attribute First, Then Generate approach tends to use more computing resources and is also slower compared to direct end-to-end generation. This is because it requires several steps, each involving a separate call to the model.
@@ Appendix A Datasets Preprocessing
For the training processes in the MDS setting, we use the alignment dataset from Ernst et al. (2022), automatically extracted from the DUC and TAC summarization datasets. [^13] https://duc.nist.gov/ For the development and test splits, we use an alignment dataset extracted from the Multi-News summarization dataset (Fabbri et al., 2019), which was assembled in the context of another work, currently under peer-review. [^14] Upon publication of that work, we will disclose its name. The dataset consists of 100 summaries with a corresponding set of source news articles, comprising an average of 3 articles per summary. Additionally, for each such pair of summaries and set of documents, the authors trained qualified annotators to identify aligning spans, following the approach delineated in Slobodkin et al. (2022).
@@ Appendix B Meta-Evaluation for AutoAIS
Figure B.7: Example input and output for the fine-tuned sentence-by-sentence generation in the MDS setting.
@@ Appendix C Implementation Details
In this section, we describe the implementation details for all of our models described in Section 3. Each component is described separately for its ICL and fine-tuning approach. RougeL and BertScore were calculated with the evaluate library  [^15] https://huggingface.co/docs/evaluate/en/index (Lin, 2004).
For our ICL experiments, we deploy the Gemini-Pro model (Gemini Team Google, 2023), last updated on December 2023, which is available via API calls. [^16] https://ai.google.dev/models/gemini In all LFQA experiments, the query is included as part of the input for every component. For all the subtasks, as well as the baselines, we perform a grid search on the optimal number of in-context examples (1-3 in the MDS setting and 1-4 in the LFQA setting) and temperature hyper-parameter (0.1,0.3,0.5,0.7,0.9) by evaluating each combination on the respective development set. For all fine-tuned experiments, we fine-tune Primera for both the MDS and LFQA settings. We also tested training LongT5, but found Primera performs significantly better. The number of parameters in the Primera model is 447M. All fine-tuned models were trained for 10 epochs and the best model checkpoint was chosen by $rougeL$, unless mentioned otherwise. The training of each model took approximately 1 hour on one Nvidia A100 80GB GPU. Hyperparameter training included 3 epochs for each model and used the following sweep configuration: learning rate (max: 5e-4; min: 5e-8), warmup steps (max: 300; min: 0), weight decay (0.0, 0.2, 0.5).

### v3

@@ Abstract
Recent efforts to address hallucinations in Large Language Models (LLMs) have focused on attributed text generation, which supplements generated texts with citations of supporting sources for post-generation fact-checking and corrections. Yet, these citations often point to entire documents or paragraphs, burdening users with extensive verification work. In this paper, we introduce a locally-attributable text generation approach, prioritizing concise attributions. Our method, named “Attribute First, then Generate”, breaks down the conventional end-to-end generation process into three intuitive steps: content selection, sentence planning, and sequential sentence generation. By initially identifying relevant source segments (“select first”) and then conditioning the generation process on them (“then generate”), we ensure these segments also act as the output’s fine-grained attributions (“select” becomes “attribute”). Tested on Multi-document Summarization and Long-form Question-answering, our method not only yields more concise citations than the baselines but also maintains—and in some cases enhances—both generation quality and attribution accuracy. Furthermore, it significantly reduces the time required for fact verification by human assessors. [^1] Our code is publicly available at https://github.com/lovodkin93/attribute-first-then-generate
@@
Figure 1: In our approach, demonstrated here on LFQA, the input to the model assumes a question and a collection of documents, which could be retrieved or provided by the user. The generated text is grounded in these documents, accompanied with sentence-level citations to concise relevant spans from the supporting sources. Previous approaches (e.g. Gao et al., 2023a, b) cite the whole source (truncated here for brevity), making it harder for readers to reliably verify the generation.
@@ 3.2 In-context Learning
For each of the substeps delineated in the previous section, we propose a prompt-based in-context learning (ICL) strategy, where a model is guided with specific instructions and carefully selected few-shot examples. The content selection step ((a) in Fig. 2) is formulated as $\mathcal{D}\rightarrow h_{1},\ldots,h_{m}$, where $\mathcal{D}$ is the set of source documents and $h_{i}$ corresponds to a verbatim span from $\mathcal{D}$. Namely, a model is tasked with outputting relevant spans from $\mathcal{D}$, separated by a designated markup (see prompts in Figs. B.1 and B.5). In MDS, these spans are aimed to capture salient information, while for LFQA, the model, presented with a query, is instructed to identify spans answering that query. These spans are then located within the source documents via string-matching, with any undetectable span being omitted (for more details, see Section C.1). Notably, in all subsequent steps, we “highlight” the identified spans within the document set $\mathcal{D}$ via designated markups before and after each span, akin to the methodology used in prior work (Gehrmann et al., 2018; Slobodkin et al., 2022, 2024).
For the sentence planning step ((b) in Fig. 2), the documents, now “highlighted”, are concatenated, truncating each at the final highlighted span to manage input length and focus the model’s attention. To highlight, we use two designated markups, <highlight_start> and <highlight_end>, which we insert into the text before and after each consecutive ‘highlighted’ span, respectively (see prompts in Figs. B.2 and B.6). The model is instructed to organize $h_{1},\ldots,h_{m}$ into an ordered set of highlight clusters $C_{1},\ldots,C_{n}$, where $C_{i}=\{h_{k}\}_{k=1}^{|C_{i}|}$, to guide the subsequent generation phase. For more details, see Section C.2.
Finally, in step (c), the model is guided to generate the output one sentence at a time, adhering to the structured plan. Specifically, we separate the generation process into a series of model calls, each focusing on a single cluster’s highlights, concatenated in the same manner as in step (b). The model is also presented with all previously generated sentences $s_{1:i}$. The task is then structured as maximizing $p(s_{i+1}|s_{1:i},C_{i+1})$, guiding the model to produce the subsequent sentence $s_{i+1}$, considering both the current cluster of highlights $C_{i+1}$ and the previously-generated sentences $s_{1:i}$. [^2] In Section 5.1 we show that conditioning on $s_{1:i}$ improves coherence. For more details, see Section C.3 as well as the full prompts structure in Figs. B.3 and B.7.
@@
We also explore a variation that combines the sentence-planning and sentence-by-sentence generation stages ((b) and (c)), inspired by the Chain-of-Thought paradigm (CoT; Wei et al., 2022). Given the selected content from step (a), the model is tasked with generating sentence-level plans, deciding on highlights to include in each synthesized sentence, alongside producing these sentences (see prompts in Figs. B.4 and B.8). The model is guided to produce the final output, after forming the plan, focusing on unifying and refining sentences for coherence. For sentence-level attribution, we link each sentence in the output with the highlights from its planning stage in the CoT process. [^3] We use spaCy sentencizer Honnibal and Montani (2017). For more details, refer to Section C.3.
@@ 3.3 Fine-tuning
We also explore fine-tuning models for the various steps of our framework, aligning the input-output formulation of each subtask to its ICL counterpart in Section 3.2. During content selection ((a) in Fig. 2), our model is taught to copy task-specific segments from $\mathcal{D}$ verbatim, and to separate them using a special markup (see Section C.1 and Figs. B.9 and B.12). To further ensure strict lexical fidelity to the source text, a constrained decoding approach is adopted during inference (see Section C.6 for more details).
In the sentence planning step ((b) in Fig. 2), we use special tokens to ‘highlight’ the previously selected spans, and teach our model to copy verbatim the “highlighted” spans in a clustered and ordered manner, using special tokens to separate between different clusters and different highlights within a cluster (see Section C.2 and Figs. B.10 and B.13 for more details). This phase similarly incorporates a constrained decoding technique to maintain exact lexical matching to the highlighted content. Finally, for the sentence generation step ((c) in Fig. 2), we use special tokens to separate between the highlighted documents, the prefix, and for the LFQA setting the query (see Section C.3 and Figs. B.11 and B.14).
@@
We also explore a joint strategy that combines the content selection and plan generation stages ((a) and (b)). Here, the entire document set $\mathcal{D}$ serves as input, and the output is an ordered series of “highlight” clusters $C_{1},\ldots,C_{n}$. To ensure exact copying of source spans, we employ the aforementioned constrained decoding strategy. For more details, see Section C.2.
@@ 4.1 Datasets
For MDS, we use an alignment dataset derived from the DUC, TAC benchmarks [^4] https://duc.nist.gov/ and the Multi-News dataset (Fabbri et al., 2019), comprising of pairs of summaries, corresponding source articles and summary-source alignments from Ernst et al. (2022, 2024). For LFQA, we rely on human-annotated alignments of large language model responses with citations Liu et al. (2023). By filtering responses that were rated as verifiable, we ensure high-quality alignment for this setting. Further details and dataset processing information are provided in Appendix A.
@@ 4.2 Models and Baselines
Table 1: MDS results of both in-context learning (ICL) and fine-tuned (FT) models. The left two columns are metrics that measure the quality of the output, ignoring attributions. The right columns report the attribution quality, including the length (number of tokens) of the cited text. No Att. shows the percentage of generated sentences lacking any attribution.
@@
Table 2: LFQA results of both in-context learning (ICL) and fine-tuned (FT) models. The left two columns are metrics that measure the quality of the output, ignoring attributions. The right columns report the attribution quality, including the length (number of tokens) of the cited text. No Att. shows the percentage of generated sentences lacking any attribution.
@@ 4.3 Evaluation
The evaluation of generated attributed texts consists of two steps. First, annotators are asked to evaluate text coherence on a 5-point Likert scale, and in the LFQA setting also its helpfulness in addressing the query. Next, annotators proceed to evaluate the quality of citations. We follow the sentence-level attribution evaluation protocol from Gao et al. (2023a), which extends the Attributable to Identified Sources (AIS; Rashkin et al., 2022) human evaluation framework of attribution to the sentence-level. Annotators judge whether each output sentence is fully supported by its attribution, scoring 1 for full support and 0 otherwise. The final AIS score is the average of these scores across sentences. For more details, see Appendix G.
@@ 5 Results and Analyses
We first note that our approach, which extracts focused spans from the input documents, leads to significantly shorter attributions than the ALCE baseline, where citations coarsely point to entire documents. For example, the Attr. FirstCoT variant’s citations are on average 45 times shorter than ALCE’s. [^8] See examples of fine-grained attributions generated by our approach in Appendix E. Importantly, we find that assessing our paradigm’s attribution takes annotators roughly half the time compared to those of ALCE, emphasizing the benefit of attribution localization in streamlining the manual fact-checking process.
In terms of attribution accuracy, our method surpasses the ALCE baseline in the LFQA domain (both on the automatic AutoAIS and the human AIS metrics), while in MDS, although ALCE shows higher accuracy, it leads to twice as many unattributed sentences. We also note that while in the MDS setting our method may decrease attribution accuracy compared to ALCE’s full-document citations, readers can always fall back to examining the full document if the provided highlights are insufficiently supporting. Indeed, analyses of the human annotations revealed that around 42% of the 109 sentences deemed unsupported were partially supported, with the remainder of the attributing details frequently found in immediate proximity to the highlighted sections (usually within the same or adjacent sentences). [^9] See partially-attributed examples in Appendix H. This observation indicates that while our sentence fusion step occasionally includes slightly more information than the selected highlights (later used as fine attributions), it mostly adds details from the surrounding context. [^10] This trend is also observed in our automatic evaluation, where extending sub-sentence attributing spans to their corresponding full sentences improves AutoAIS. See Appendix D. Hence, while suggesting room for improvement in strict adherence to highlights, this does not significantly hinder fact-checking effectiveness.
@@ 5.1 Do Models Benefit from Seeing the Prefix?
We also examine whether providing models with the context of previously generated sentences indeed aids the generation process, particularly in enhancing output cohesion as defined by Maimon and Tsarfaty (2023). Cohesion involves sentences being “referentially linked”, e.g., via co-referring mentions, or linked through discourse connectors. To analyze the impact of the prefix on cohesion, we modified the sentence generation prompt to exclude the prefix and conducted the sentence fusion step (step (c) in Fig. 2) using the ICL variant for both MDS and LFQA tasks, maintaining identical sentence-level plans. Starting from the second sentence to isolate the impact of the prefix, we examined 20 instances per task. Our analysis of 113 sentences reveals that approximately 30% were more cohesive when including the prefix. Specifically, 23% showed improved referential linkage, mostly attributed to better use of co-reference, and 7% demonstrated better discourse connector usage. [^11] See Appendix I for examples. Additionally, our findings suggest that prefixes play an increasing role in cohesion as the output progresses (see Fig. 3), implying that cohesion gains importance as the text develops. [^12] The observed plateau in cohesion levels in LFQA is attributed to its outputs generally not exceeding three sentences.
@@ 7 Conclusion
Our Attribute First, Then Generate approach tends to use more computing resources and is also slower compared to direct end-to-end generation, as detailed in Appendix J. This is because it requires several steps, each involving a separate call to the model.
@@
Acknowledgements
This work was supported by the Israel Science Foundation (grant no. 2827/21).
@@ Appendix A Datasets Preprocessing
For the training processes in the MDS setting, we use the alignment dataset from Ernst et al. (2022), automatically extracted from the DUC and TAC summarization datasets. [^14] https://duc.nist.gov/ For the development and test splits, we use the alignment dataset from Ernst et al. (2024), extracted from the Multi-News summarization dataset (Fabbri et al., 2019). The dataset consists of 100 summaries with a corresponding set of source news articles, comprising an average of 3 articles per summary. Additionally, for each such pair of summaries and set of documents, the authors trained qualified annotators to identify aligning spans, following the approach delineated in Slobodkin et al. (2022).
@@ Appendix B Meta-Evaluation for AutoAIS
Figure B.5: Example prompt for the content selection subtask in the LFQA setting.
Figure B.6: Example prompt for the clustering subtask in the LFQA setting.
Figure B.7: Example prompt for iterative, sentence-by-sentence generation approach in the LFQA setting.
Figure B.8: Example prompt for the CoT approach in the LFQA setting.
@@
Figure B.11: Example input and output for the fine-tuned sentence-by-sentence generation in the MDS setting. This is the first sentence generated in the summary, so there is no prefix.
Figure B.12: Example input and output for the fine-tuned content selection model in the LFQA setting. Highlight start and end markups are colored blue.
Figure B.13: Example input and output for the fine-tuned sentence planning model in the LFQA setting. Highlights separators are colored blue and cluster separators are colored green.
Figure B.14: Example input and output for the fine-tuned sentence-by-sentence generation in the LFQA setting. This is the first sentence generated in the summary, so there is no prefix.
@@ Appendix C Implementation Details
In this section, we describe the implementation details for all the component described in Section 3, both the ICL and fine-tuning approaches. RougeL and BertScore were calculated with the evaluate library  [^15] https://huggingface.co/docs/evaluate/en/index (Lin, 2004).
For our ICL experiments, we use the Gemini-Pro model (Gemini Team Google, 2023), last updated on December 2023, which is available via API calls. [^16] https://ai.google.dev/models/gemini In all LFQA experiments, the query is included as part of the input for each component. For all subtasks and baselines, we perform a grid search on the optimal number of in-context examples (1-3 in the MDS setting and 1-4 in the LFQA setting) and temperature hyperparameter (0.1,0.3,0.5,0.7,0.9) by evaluating each combination on the respective development set. For all fine-tuned variants, we fine-tune Primera for both tasks. We also tested training LongT5, but found Primera performs significantly better. The number of parameters in the Primera model is 447M. All fine-tuned models were trained for 10 epochs and the best model checkpoint was chosen by RougeL, unless mentioned otherwise. The training of each model took approximately 1 hour on one Nvidia A100 80GB GPU. Hyperparameter training included 3 epochs for each model and used the following sweep configuration: learning rate (max: 5e-4; min: 5e-8), warmup steps (max: 300; min: 0), weight decay (0.0, 0.2, 0.5).